Please help me solve this word problem: Mark is 11 years older that his sister. In 8 years he will be twice as old as she will be then. How old are they now?
Here's what I have so far: m=s+11+2s+8. Please help.

Answer by nerdybill(5404)

Mark is 11 years older that his sister. In 8 years he will be twice as old as she will be then. How old are they now?
Here's what I have so far: m=s+11+2s+8.
.
Let x = age of sister
then because "Mark is 11 years older that his sister."
x+11 = age of Mark
.
From: "In 8 years he will be twice as old as she will be then."
we get our equation of:
(x+11) + 8 = 2(x+8)
x + 11 + 8 = 2x + 16
x + 19 = 2x + 16
x + 3 = 2x
3 years = x (Mark's sister's age)
.
Mark's age:
x+11 = 3+11 = 14 years (Mark's age)